OpenStudy (anonymous):

A 8 kg box is placed on a mass balance. A string exerts a force of 22.8 N upwards on the box. What does the mass balance read (in kg)?

OpenStudy (doc.brown):

\[F=ma\]\[a_{Earth}=g=-9.8m/s^2\]\[F=mg\]\[\sum F=mg+22.8N\]\[F=8kg(-9.8m/s^2)+22.8N\]\[F=-55.6N\]\[m=\frac{F}{a}\]\[m=\frac{-55.6kg\cdot m/s^2}{-9.8m/s^2}\]\[m=\]